OpenStudy (anonymous):

ln 3x^2 over (x+1)^10 what its expand?

OpenStudy (jamesj):

\[\ln \left( \frac{3x^2}{(x+1)^{10}} \right)\] yes?

OpenStudy (anonymous):

yes thats the statement

OpenStudy (jamesj):

This is equal to =ln(3x^2) - ln[ (x+1)^10 ] =ln 3 + ln(x^2) - 10.ln(x+1) =ln 3 + 2.ln x - 10.ln(x+1)

OpenStudy (jamesj):

because A) ln(ab) = ln a + ln b B) ln(a^b) = b.ln a

OpenStudy (jamesj):

from B) it follows also that C) ln(1/b) = - ln b and from A) and B) D) ln(a/b) = ln a - ln b I used all four A)-D) above

OpenStudy (anonymous):

thank you :)